### v1

@@ 5.1.1 Mathematical formulation and architecture choices
that maps samples $z$ from a latent distribution $\rho$—typically a standard isotropic Gaussian—to the parameter space. That is,
@@
Maximum likelihood and invertible transport maps
When $\mathsf{T}_{\theta}$ is invertible (e.g., a normalizing flow), the density of the prior $\mu_{\theta}$ is
$\mu_{\theta}(u)=\rho\bigl{(}\mathsf{T}^{-1}_{\theta}(u)\bigr{)}\big{\lvert}\det\bigl{(}\nabla\mathsf{T}^{-1}_{\theta}(u)\bigr{)}\big{\rvert}\,$. (5.9)
The parameters $\theta\in\Theta$ can be optimized using maximum likelihood estimation
$\max_{\theta\in\Theta}\frac{1}{N}\sum_{n=1}^{N}\log\mu_{\theta}(u_{n})$ (5.10)
by plugging in the density formula (5.9).
@@
Architectures based on autoencoders consist of an encoder-decoder pair
@@
that are trained to reconstruct input data. This amounts to the optimization problem
@@
After training, the prior over the space $\mathcal{U}$ is implicitly defined by pushing forward a simple distribution $\rho\in\mathscr{P}(\mathcal{Z})$. The decoder $\mathsf{T}_{\theta}$ maps this latent distribution to the parameter space. This induces a trained prior distribution $\mu_{\theta}=(\mathsf{T}_{\theta})_{\#}\rho$ as before.
Evidence lower bound and variational autoencoders
Variational autoencoders (VAEs) are generative models that represent the prior by introducing a latent random variable $z\in\mathcal{Z}$, which is typically drawn from a standard Gaussian $\rho=\mathcal{N}(0,I)$ [260, 261]. The conditional likelihood of the parameter $u$ given a latent variable $z$ is modeled as
$u\sim p_{\theta}({\,\cdot\,}\,|\,z)\,.$ (5.13)
The probability measure $p_{\theta}({\,\cdot\,}\,|\,z)$ is parametrized by $z\in\mathcal{Z}$ and is known as a decoder network. It is common to take
$p_{\theta}({\,\cdot\,}\mid z)=\mathcal{N}\bigl{(}m_{\theta}(z),\Sigma_{\theta}(z)\bigr{)}$ (5.14)
to be a Gaussian distribution. The mean $m_{\theta}\colon\mathcal{Z}\to\mathcal{U}$ and covariance $\Sigma_{\theta}\colon\mathcal{Z}\to\mathcal{L}(\mathcal{U};\mathcal{U})$ are parametrized by neural networks or neural operators. The covariance is usually constrained to be diagonal.
Since the marginal likelihood
$p_{\theta}(du)\coloneqq\operatorname{\mathbb{E}}_{z\sim\rho}\bigl{[}p_{\theta}(du\,|\,z)\bigr{]}$ (5.15)
can be intractable to compute, we approximate the posterior $p_{\theta}(dz\,|\,u)$ with a variational distribution $q_{\vartheta}(dz\,|\,u)$. This is often a Gaussian encoder
$q_{\vartheta}({\,\cdot\,}\,|\,u)=\mathcal{N}\bigl{(}\widetilde{m}_{\vartheta}(u),\widetilde{\Sigma}_{\vartheta}(u)\bigr{)}\,.$ (5.16)
Just as for the decoder, the mean $\widetilde{m}_{\vartheta}\colon\mathcal{U}\to\mathcal{Z}$ and covariance $\widetilde{\Sigma}_{\vartheta}\colon\mathcal{U}\to\mathcal{L}(\mathcal{Z};\mathcal{Z})$ of the encoder are parametrized by neural networks or neural operators. The covariance is typically enforced to be diagonal.
VAE training maximizes the evidence lower bound (ELBO). The ELBO is a tractable lower bound on $\log p_{\theta}(u)$, given by
@@
$\begin{split}\log p_{\theta}(u)&\geq\mathrm{ELBO}(\theta,\vartheta;u)\\ &\coloneqq\mathbb{E}_{z\sim q_{\vartheta}({\,\cdot\,}\,|\,u)}\bigl{[}\log p_{\theta}(u\,|\,z)\bigr{]}-\mathsf{KL}\big{(}q_{\vartheta}({\,\cdot\,}\,|\,u)\;\big{\|}\;\rho\big{)}\,.\end{split}$ (5.17)
The first term of the ELBO encourages accurate reconstruction of $u$. The second term regularizes the encoder by keeping it close to the latent prior distribution $\rho$. The mean of the decoder, $m_{\theta}$, can be used as a generator. This means that after training, samples from the prior $z\sim\rho$ are mapped to the parameter space with $u=m_{\theta}(z)$. This induces a learned prior distribution $\mu_{\theta}\coloneqq(m_{\theta})_{\#}\rho$ over $u$.
@@
where $\mu_{\mathrm{data}}$ is the training data distribution (5.5). The discriminator tries to assign $1$ to real data samples and 0 to generated ones, while the generator attempts to fool the discriminator by producing samples for which $D_{\vartheta}(\mathsf{T}_{\theta}(z))\approx 1$. Variants such as the Wasserstein GAN [262] replace the Jensen–Shannon divergence in the GAN loss with an approximation of the dual form of the Wasserstein-$1$ distance. This leads to the minimax game
$\min_{\theta}\max_{{\mathopen{}\mathclose{{\left\{\vartheta\,\middle|\,\operatorname{Lip}(D_{\vartheta})\leq 1}}\right\}}}\Bigl{\{}\mathbb{E}_{u\sim\mu_{\mathrm{data}}}\bigl{[}D_{\vartheta}(u)\bigr{]}-\mathbb{E}_{z\sim\rho}\bigl{[}D_{\vartheta}\bigl{(}\mathsf{T}_{\theta}(z)\bigr{)}\bigr{]}\Bigr{\}}\,.$ (5.22)
@@
Finally, some methods seek to learn an explicit, parametrized regularization functional $R_{\theta}\colon\mathcal{U}\to\mathbb{R}$ that approximates the negative log prior density from (5.3). In approaches based on maximum likelihood estimation, one fits a model of the form
$\mu_{\theta}(u)\propto\exp\bigl{(}-R_{\theta}(u)\bigr{)}$ (5.23)
to a dataset $\{u_{n}\}_{n=1}^{N}\sim\mu_{\mathrm{data}}^{\otimes N}$ by solving
$\max_{\theta\in\Theta}\frac{1}{N}\sum_{n=1}^{N}\log\mu_{\theta}(u_{n})=\min_{\theta}\frac{1}{N}\sum_{n=1}^{N}\Bigl{[}R_{\theta}(u_{n})+\log Z_{\theta}\Bigr{]}\,$, (5.24)
where $Z_{\theta}$ is the normalization constant implied in (5.23). In practice, computing $Z_{\theta}$ is often intractable, which necessitates approximate training schemes such as contrastive divergence [263] or score matching [264]. Once trained, the resulting $R_{\theta}$ provides an explicit regularizer suitable for use in downstream inverse problem solvers.
Alternatively, score-based models, such as denoising score matching and diffusion models [36, 254], bypass the need to evaluate or approximate the normalizing constant $Z_{\theta}$ by instead learning the score function
@@
directly with neural networks or neural operators [239, 240]. While these models do not yield an explicit density for $\mu$, they enable sampling and approximate inference through stochastic dynamics. Consequently, they define the prior implicitly through its score or generative process rather than an explicit functional form.
@@ 5.1.2 Theoretical results
The quantitative robustness of Bayesian inference has been a subject of extensive study for several decades [266, 267, 268, 269]. Recent advancements in the theoretical analysis of Bayesian inverse problems and generative modeling have reignited interest in this topic [270, 271, 272, 273]. A key question of relevance to this chapter is the robustness of the posterior distribution with respect to perturbations in the prior distribution. Such prior perturbations can arise from randomized approximations [274, 176] or finite element approximations [275] of Gaussian process priors, for example. To present the results, we adopt the compact linear functional notation $\pi(f)\coloneqq\operatorname{\mathbb{E}}_{u\sim\pi}[f(u)]$, which represents the integral of a function $f$ with respect to a measure $\pi$.
@@
for any $p\in[1,\infty]$; see [272, Lem. 5.1, p. 38]. The constant $C_{p}(\mu,\nu;y)$ depends on the p/(p1)p/(p-1)-th moment of the priors and is known explicitly [272]. In particular, the limit $p\to 1$ requires one of the priors to have compact support, which excludes popular unbounded priors such as Gaussian measures. As a result, the scenario $\mathsf{d}=\mathsf{d}_{0}=\mathsf{W}_{1}$ is relatively uncommon.
@@
The bounded Lipschitz distance is an integral probability metric defined as
$\mathsf{d}_{\mathrm{BL}}(\mu,\nu)\coloneqq\sup_{\lVert f\rVert_{\mathrm{BL}}\leq 1}\lvert\mu(f)-\nu(f)\rvert\,,\ \text{ where }\ \lVert f\rVert_{\mathrm{BL}}\coloneqq\max\bigl{(}\lVert f\rVert_{L^{\infty}},\operatorname{Lip}(f)\bigr{)}$ (5.28)
for any probability measures $\mu$ and $\nu$. We recall that $\operatorname{Lip}(f)=\sup_{u\neq v}\lvert f(u)-f(v)\rvert/\lVert u-v\rVert$. The distance $\mathsf{d}_{\mathrm{BL}}$ metrizes the topology of weak convergence [80]. Local Lipschitz stability in Dudley’s metric has been established in previous work [266, 267]. However, we present a more quantitative version tailored for inverse problems that is inspired by prior work [272, Remark 5.2, p. 39].
@@
We provide the proof of Proposition 1 below, which is prototypical of how other basic results of the prior-to-posterior type (5.26) are established (see, e.g., [34, Thm. 4.6, p. 43] or [270, Thm. 8, p. 11]).
@@
Weak metrics such as $\mathsf{d}_{\mathrm{BL}}$ and $\mathsf{W}_{p}$ are particularly well-suited for data-driven prior learning because these distances remain meaningful even when the two input distributions are mutually singular. In contrast, statistical divergences such as TV, KL, and Hellinger attain their maximum value when the measures are mutually singular. This issue often arises in practice because a finite training dataset can lead to an approximate prior that is mutually singular with respect to the reference prior. For instance, using the empirical measure as an estimator results in discrete mismatched support, while employing the closest Gaussian approximation to the reference prior yields a degenerate covariance operator. More fundamentally, in infinite-dimensional spaces, probability measures tend to be mutually singular. As an illustrative example, we refer to the Feldman–Hajek equivalence theorem for Gaussian measures [57, Thm. 6.13, p. 531].
@@
Theoretical analyses of regularizer learning for linear inverse problems in infinite dimensions have recently gained attention, with notable developments presented in [277, 278, 138, 279, 280]. This line of research is more naturally situated within the nonparametric statistical setting described in Section 2.1.3, as opposed to the functional-analytic viewpoint of Section 2.2 or the Bayesian perspective in Section 2.1.1. However, the research in this direction can still be related to Bayesian inference by interpreting variational point estimators as either the mean or MAP point of an underlying posterior distribution. Such a connection will be made clear in the following discussions. Regardless of viewpoint, the probabilistic models for the parameters and observed data in the inverse problem play a major role in the theoretical insights that are unveiled, as we now explain.
@@
The element $h\in\mathcal{U}$ is a centering parameter for $u$. The continuous linear operator $B\in\mathcal{L}(\mathcal{U};\mathcal{U})$ penalizes the smoothness of $u$. We interpret the second term on the right-hand side of (5.32) as being equal to $\infty$ if $(u-h)\notin\operatorname{Ran}(B)$. The first term can be interpreted formally as $\lVert\Gamma^{-1/2}(y-\mathcal{A}u)\rVert_{\mathcal{Y}}^{2}$, but such an expression is infinite almost surely if $\mathcal{Y}$ is infinite-dimensional. To address this, the actual discrepancy function $\mathsf{d}_{\Gamma}$ captures the technical modifications needed to ensure that (5.32) is well-defined [277, Sec. 2]. We emphasize that $\mathsf{d}_{\Gamma}$ depends on the covariance operator $\Gamma$ of the noise, which is assumed known. Thus, $\mathcal{R}_{h,B}^{\Gamma}$ itself depends on the inherent noise level in the observations.
Due to the possibility of a nonzero $h$, the map $\mathcal{R}_{h,B}^{\Gamma}$ is affine rather than linear. Its closed-form expression is provided in [277, Sec. 2, p. 5]. The associated calculations are similar to those used to establish that the posterior distribution for a linear Gaussian Bayesian inverse problem in an infinite-dimensional setting (for both data and parameters) is well-defined [67]. However, it is important to emphasize that no Gaussian assumptions are made in the current context.
@@
The expectation is taken with respect to the joint distribution $\nu\coloneqq\operatorname{Law}(u,y)$ of the random variable $(u,y)$ under (5.30). Under minor technical assumptions, the unique global minimizer of (5.33) is $(h^{\star},B^{\star})=(\overline{u},\Sigma^{1/2})$ [277, Thm. 3.1, p. 6]. Notably, even though (5.33) depends on the full joint distribution $\nu$—and hence the forward map $\mathcal{A}$—the optimal solution only depends on the first two moments of the marginal law of $u$. The technical contributions leading to this result are substantial, as they require careful analysis to handle the infinite-dimensional noise process $\eta$ and regularized inverses of compact operators.
It is also possible to develop finite-sample estimators for the problem (5.33) if only a dataset of i.i.d. pairs $\{(u_{n},y_{n})\}_{n=1}^{N}$ from $\nu$ is available. Because the optimal Tikhonov regularizer $(h^{\star},B^{\star})$ is known to be the mean and square root covariance of the random variable $u$, a natural unsupervised strategy is first to compute the empirical mean $\widehat{u}$ and empirical covariance $\widehat{\Sigma}$ from the training data, then define the plug-in estimator $(\widehat{h},\widehat{B})\coloneqq(\widehat{u},\widehat{\Sigma}^{1/2})$. Assuming that $u$ is strongly-subgaussian, application of tools from covariance estimation [277, Thm. 4.2, p. 8, and p. 29] deliver the nonasymptotic error bound

### v2

@@ 5.1.1 Mathematical formulation and architecture choices
that maps samples $z$ from a latent distribution $\rho$—typically a standard isotropic Gaussian—to the parameter space. That is, the generative model
@@
Maximum likelihood and invertible transport maps.
When $\mathsf{T}_{\theta}$ is invertible (e.g., a normalizing flow), the density of the induced prior $\mu_{\theta}$ can be obtained via the change-of-variables formula as
$\mu_{\theta}(u)=\rho\bigl(\mathsf{T}^{-1}_{\theta}(u)\bigr)\big\lvert\det\bigl(\nabla\mathsf{T}^{-1}_{\theta}(u)\bigr)\big\rvert\,$, (5.9)
where $\nabla\mathsf{T}_{\theta}^{-1}$ denotes the Jacobian matrix of the inverse transformation. The parameters $\theta\in\Theta$ can then be optimized using maximum likelihood estimation,
$\max_{\theta\in\Theta}\frac{1}{N}\sum_{n=1}^{N}\log\mu_{\theta}(u_{n})\,$, (5.10)
by substituting in the density expression (5.9).
Normalizing flows (NFs) provide a flexible class of invertible parametrizations $\mathsf{T}_{\theta}$ that make the optimization problem (5.10) tractable by exploiting the triangular or auto-regressive structure of the Jacobian determinant; see, for example, [288, 289, 290]. Related continuous-time generalizations, known as neural ordinary differential equation (ODE) flows, define $\mathsf{T}_{\theta}$ as the finite time flow map of a learned ODE and interpret the log-determinant term through the instantaneous trace of the Jacobian [291, 292].
The preceding constructions of $\mathsf{T}_{\theta}$, however, rely fundamentally on the finite-dimensional change-of-variables formula (5.9) and Lebesgue probability densities on $\mathbb{R}^{d}$. Extending such density-based techniques to infinite-dimensional settings, such as function spaces arising in PDE-based inverse problems, is nontrivial and restrictive because Gaussian reference measures on Banach or Hilbert spaces are typically mutually singular [61]. Nevertheless, there are promising recent efforts to generalize normalizing flows to function space [264]. This functional normalizing flow approach aims to transport Gaussian process priors through invertible transformations while maintaining measure-theoretic rigor, potentially bridging finite-dimensional NF methods with infinite-dimensional Bayesian inverse problem formulations. See also the discussion in Section 4.4.
@@
Architectures based on autoencoders consist of an encoder–decoder pair
@@
trained jointly to reconstruct input data. This leads to the optimization problem
@@
Depending on the setting, $(E_{\vartheta},\mathsf{T}_{\theta})$ may represent a FtV encoder and a VtF decoder pair (recall Section 4.2.1), or remain entirely finite-dimensional. In either case, the formulation (5.12) only relies on the Banach space norm on $\mathcal{U}$, making it applicable to both discrete and functional representations. One can also consider regularizing the encoded latent variables in the objective function. This naturally motivates a Gaussian reference $\rho\in\mathscr{P}(\mathcal{Z})$ that is consistent with the role of latent distributions in functional autoencoders [190]. After training, this reference is pushed forward through the decoder to define the induced distribution $\mu_{\theta}=(\mathsf{T}_{\theta})_{\#}\rho$, which serves as a learned prior generative model over the space $\mathcal{U}$.
Variational autoencoders
Variational autoencoders (VAEs) extend the deterministic framework of autoencoders by introducing a probabilistic generative model. While a standard autoencoder learns deterministic encoder and decoder mappings $E_{\vartheta}\colon\mathcal{U}\to\mathcal{Z}$ and $\mathsf{T}_{\theta}\colon\mathcal{Z}\to\mathcal{U}$ that minimize a reconstruction loss, e.g., (5.12), a VAE treats both mappings as stochastic maps taking value in spaces of probability measures, i.e., Markov kernels. In particular, the encoder outputs a distribution $q_{\vartheta}({\,\cdot\,}\mid u)$ over latent variables. The decoder specifies a conditional distribution $p_{\theta}({\,\cdot\,}\mid z)$ over reconstructions.
VAEs can be formulated from a joint-measure and transport perspective, which is compatible with both finite- and infinite-dimensional settings [36, Sec. 13.6]. In this formulation, one defines a joint probability measure
$p_{\theta}(dz,du)=p_{\theta}(du\mid z)\,\rho(dz)$ (5.13)
on the product space $\mathcal{Z}\times\mathcal{U}$. Here, $\rho$ is a fixed reference measure on the latent space $\mathcal{Z}$ (typically $\rho=\mathcal{N}(0,I)$ in finite dimensions) and $p_{\theta}({\,\cdot\,}\mid z)$ is a parametrized decoder operator that maps each latent $z$ to a probability measure on $\mathcal{U}$. This construction induces a stochastic generative model
$z\sim\rho\quad\text{and}\quad u\sim p_{\theta}({\,\cdot\,}\mid z)$ (5.14)
that specifies the distribution of $u$ through the composition of a latent draw and a conditional sampling step.
In practice, the conditional measures $p_{\theta}({\,\cdot\,}\mid z)$ are often taken to be Gaussian,
$p_{\theta}({\,\cdot\,}\mid z)=\mathcal{N}\bigl(m_{\theta}(z),\Sigma_{\theta}(z)\bigr)\,$, (5.15)
where the mean $m_{\theta}\colon\mathcal{Z}\to\mathcal{U}$ and covariance operator $\Sigma_{\theta}\colon\mathcal{Z}\to\mathcal{L}(\mathcal{U};\mathcal{U})$ are represented by neural networks or neural operators. This viewpoint interprets the decoder as a (possibly stochastic) operator that maps latent representations to probability measures over functions or parameters, enabling discretization-invariant modeling of infinite-dimensional data.
To infer latent representations from data, one introduces an encoder operator $q_{\vartheta}({\,\cdot\,}\mid u)$ mapping $u\in\mathcal{U}$ to distributions on $\mathcal{Z}$ and defines a corresponding variational joint measure
$q_{\vartheta}(dz,du)=q_{\vartheta}(dz\mid u)\,\nu(du)\,$, (5.16)
where $\nu$ denotes the (empirical version of the) data measure $\mu_{\mathrm{data}}$ on $\mathcal{U}$. The encoder is also often taken to be Gaussian in practice.
The goal of VAE training is to bring the model joint $p_{\theta}(dz,du)$ and the variational joint $q_{\vartheta}(dz,du)$ into alignment. This can be achieved by minimizing a divergence such as
$\mathsf{KL}(q_{\vartheta}\;\|\;p_{\theta})$ (5.17)
over $(\vartheta,\theta)$, which compares the two joint measures directly. The minimization implicitly balances reconstruction of $u$ through the decoder operator and regularization of the encoder against the latent reference measure $\rho$. Recent work such as [190] instates the joint-measure-based VAE framework in infinite-dimensional function spaces. However, such infinite-dimensional VAEs necessitate strong assumptions on the data distribution $\mu_{\mathrm{data}}$ in order for the KL divergence between the joint measures to be finite [190, Secs. 2–3]. Samples from the trained VAE may be obtained through (5.14).
In finite-dimensional settings, training a VAE can be interpreted as maximizing the evidence lower bound (ELBO), which provides a tractable lower bound on the marginal log-likelihood $\log p_{\theta}(u)$. Here $p_{\theta}(u)$ denotes the model evidence, which is given by marginalizing over the latent variable via
$p_{\theta}(du)\coloneqq\operatorname{\mathbb{E}}_{z\sim\rho}\bigl[p_{\theta}(du\mid z)\bigr]\,.$ (5.18)
The ELBO takes the form
@@
$\begin{split}\log p_{\theta}(u)&\geq\mathrm{ELBO}(\theta,\vartheta;u)\\ &\coloneqq\operatorname{\mathbb{E}}_{z\sim q_{\vartheta}({\,\cdot\,}\mid u)}\bigl[\log p_{\theta}(u\mid z)\bigr]-\mathsf{KL}\big(q_{\vartheta}({\,\cdot\,}\mid u)\;\big\|\;\rho\big)\,.\end{split}$ (5.19)
The first term encourages the decoder to reconstruct $u$ accurately from samples of the latent variable $z$, while the second term regularizes the encoder by keeping $q_{\vartheta}({\,\cdot\,}\mid u)$ close to the reference distribution $\rho$. After minimizing the ELBO, (5.14) delivers samples from the trained VAE.
@@
where $\mu_{\mathrm{data}}$ is the training data distribution (5.5). The objective (5.23) seeks to minimize the Jensen–Shannon (JS) divergence between the model distribution $\mu_{\theta}=(\mathsf{T}_{\theta})_{\#}\rho$ and the data distribution $\mu_{\mathrm{data}}$ [293]. The discriminator tries to assign $1$ to real data samples and 0 to generated ones, while the generator attempts to fool the discriminator by producing samples for which $D_{\vartheta}(\mathsf{T}_{\theta}(z))\approx 1$. Variants such as the Wasserstein GAN [294] replace the Jensen–Shannon divergence in the GAN loss with an approximation of the dual form of the Wasserstein-$1$ distance. This leads to the minimax game
$\min_{\theta}\max_{\vartheta\in{\mathopen{}\mathclose{{\left\{{\vartheta^{\prime}}\,\middle|\,\operatorname{Lip}(D_{{\vartheta^{\prime}}})\leq 1}}\right\}}}\Bigl\{\mathbb{E}_{u\sim\mu_{\mathrm{data}}}\bigl[D_{\vartheta}(u)\bigr]-\mathbb{E}_{z\sim\rho}\bigl[D_{\vartheta}\bigl(\mathsf{T}_{\theta}(z)\bigr)\bigr]\Bigr\}\,.$ (5.24)
@@
Recent work has extended GANs to infinite-dimensional setups. In [295], Wasserstein GANs were extended to function spaces through the Generative Adversarial Neural Operator (GANO), in which the generator and discriminators are formulated as neural mappings [82]: the generator acts as a VtF map producing samples $u\in\mathcal{U}$, while the discriminator serves as a FtV map evaluating functions to scalar scores. Training uses a Wasserstein-$1$ objective with a discretization-invariant Lipschitz constraint. A complementary theoretical framework for adversarial learning on infinite-dimensional Banach spaces was provided in [296], showing convergence of the JS divergence under smoothness and positivity assumptions on the relevant Radon–Nikodym derivatives. A related work on Wasserstein GANs in Banach space is [297].
@@
Finally, some methods seek to learn an explicit, parametrized regularization functional $R_{\theta}\colon\mathcal{U}\to\mathbb{R}$ that approximates the negative log prior density from (5.3). Models of this type are often referred to as energy-based models (EBMs), since they represent probability measures through an unnormalized energy function:
$\mu_{\theta}(u)\propto\exp\bigl(-R_{\theta}(u)\bigr)\,.$ (5.25)
In approaches based on maximum likelihood estimation, one fits a density model (5.25) to a dataset $\{u_{n}\}_{n=1}^{N}\sim\mu_{\mathrm{data}}^{\otimes N}$ by solving
$\max_{\theta\in\Theta}\frac{1}{N}\sum_{n=1}^{N}\log\mu_{\theta}(u_{n})=\min_{\theta\in\Theta}\frac{1}{N}\sum_{n=1}^{N}\Bigl[R_{\theta}(u_{n})+\log Z_{\theta}\Bigr]\,$, (5.26)
where $Z_{\theta}$ is the normalization constant implied in (5.25). The only difference from (5.9) and (5.10) is the parametrization of the Lebesgue density of $\mu_{\theta}$. In practice, computing $Z_{\theta}$ is often intractable, which necessitates approximate training schemes such as contrastive divergence [298] or score matching [299]. Once trained, the resulting $R_{\theta}$ provides an explicit regularizer suitable for use in downstream inverse problem solvers. Although the Lebesgue density parametrization (5.25) is only valid in finite dimensions, it is easy to generalize to infinite dimensions by defining the density with respect to a reference probability measure $\rho$ such as a proper Gaussian measure. This leads to
$\mu_{\theta}(du)\propto\exp\bigl(-R_{\theta}(u)\bigr)\rho(du)$ (5.27)
and an analogous optimization problem to (5.26).
Alternatively, score-based models—such as denoising score matching and diffusion models [38, 282]—bypass the need to evaluate or approximate the normalizing constant $Z_{\theta}$ by instead learning the score function
@@
directly with neural networks or neural operators [262, 300, 263]. We stress that defining the score is considerably more delicate in function space settings because probability measures on infinite-dimensional Banach spaces lack a Lebesgue density. One must instead identify a generalization of the score that remains well defined without relying on the log-density gradient formula (5.28). In [301], the score is formulated explicitly in function space as the logarithmic derivative of a perturbed measure in Cameron–Martin space directions. Sampling is performed via function-valued Langevin dynamics. Alternatively, [302] introduces a conditional expectation representation that plays the role of the score. Both infinite-dimensional score formulas agree with (5.28) if the Banach space is $\mathbb{R}^{d}$.
While score-based models do not yield an explicit density for $\mu$, they enable sampling and approximate inference through stochastic dynamics. Consequently, they define a learned prior implicitly through its score or generative process rather than an explicit functional form.
@@ 5.1.2 Theoretical results
The quantitative robustness of Bayesian inference has been a subject of extensive study for several decades [304, 305, 306, 307]. Recent advancements in the theoretical analysis of Bayesian inverse problems and generative modeling have reignited interest in this topic [308, 309, 310, 311]. A key question of relevance to this chapter is the robustness of the posterior distribution with respect to perturbations in the prior distribution. Such prior perturbations can arise from randomized approximations [312, 186] or finite element approximations [313] of Gaussian process priors, for example. We now overview results of this type.
@@
for any $p\in[1,\infty]$; see [310, Lem. 5.1, p. 38]. The constant $C_{p}(\mu,\nu;y)$ depends on the p/(p1)p/(p-1)-th moment of the priors and is known explicitly [310]. In particular, the limit $p\to 1$ requires one of the priors to have compact support, which excludes popular unbounded priors such as Gaussian measures. As a result, the scenario $\mathsf{d}=\mathsf{d}_{0}=\mathsf{W}_{1}$ is relatively uncommon.
@@
We now adopt the compact linear functional notation $\pi(f)\coloneqq\operatorname{\mathbb{E}}_{u\sim\pi}[f(u)]$, which represents the integral of a function $f$ with respect to a measure $\pi$. The bounded Lipschitz distance is an integral probability metric defined as
$\mathsf{d}_{\mathrm{BL}}(\mu,\nu)\coloneqq\sup_{\lVert f\rVert_{\mathrm{BL}}\leq 1}\lvert\mu(f)-\nu(f)\rvert\,,\ \text{ where }\ \lVert f\rVert_{\mathrm{BL}}\coloneqq\max\bigl(\lVert f\rVert_{L^{\infty}},\operatorname{Lip}(f)\bigr),$ (5.31)
for any probability measures $\mu$ and $\nu$. We recall that $\operatorname{Lip}(f)=\sup_{u\neq v}\lvert f(u)-f(v)\rvert/\lVert u-v\rVert$. The distance $\mathsf{d}_{\mathrm{BL}}$ metrizes the topology of weak convergence [84]. Local Lipschitz stability in Dudley’s metric has been established in previous work [304, 305]. However, we present a more quantitative version tailored for inverse problems that is inspired by prior work [310, Remark 5.2, p. 39].
@@
We provide the proof of Proposition 1 below, which is prototypical of how other basic results of the prior-to-posterior type (5.29) are established (see, e.g., [36, Thm. 4.6, p. 43] or [308, Thm. 8, p. 11]).
@@
Weak metrics such as $\mathsf{d}_{\mathrm{BL}}$ and $\mathsf{W}_{p}$ are particularly well-suited for data-driven prior learning because these distances remain meaningful even when the two input distributions are mutually singular. In contrast, statistical divergences such as TV, KL, and Hellinger attain their maximum value when the measures are mutually singular. This issue often arises in practice because a finite training dataset can lead to an approximate prior that is mutually singular with respect to the reference prior. For instance, using the empirical measure as an estimator results in discrete mismatched support, while employing the closest Gaussian approximation to the reference prior yields a degenerate covariance operator. More fundamentally, in infinite-dimensional spaces, probability measures tend to be mutually singular. As an illustrative example, we refer to the Feldman–Hajek equivalence theorem for Gaussian measures [61, Thm. 6.13, p. 531].
@@
Theoretical analyses of regularizer learning for linear inverse problems in infinite dimensions have recently gained attention, with notable developments presented in [315, 316, 148, 317, 318]. This line of research is more naturally situated within the nonparametric statistical setting described in Section 2.1.3, as opposed to the functional-analytic viewpoint of Section 2.2 or the nonparametric Bayesian perspective in Section 2.1.1. However, the research in this direction can still be related to Bayesian inference by interpreting variational point estimators as either the mean or MAP point of an underlying posterior distribution. Such a connection will be made clear in the following discussions. Regardless of viewpoint, the probabilistic models for the parameters and observed data in the inverse problem play a major role in the theoretical insights that are unveiled, as we now explain.
@@
The element $h\in\mathcal{U}$ is a centering parameter for $u$. The continuous linear operator $B\in\mathcal{L}(\mathcal{U};\mathcal{U})$ penalizes the smoothness of $u$. We interpret the second term on the right-hand side of (5.35) as being equal to $\infty$ if $(u-h)\notin\operatorname{Ran}(B)$. The first term can be interpreted formally as $\lVert\Gamma^{-1/2}(y-\mathcal{A}u)\rVert_{\mathcal{Y}}^{2}$, but such an expression is infinite almost surely if $\mathcal{Y}$ is infinite-dimensional. To address this, the actual discrepancy function $\mathsf{d}_{\Gamma}$ captures the technical modifications needed to ensure that (5.35) is well-defined [315, Sec. 2]. We emphasize that $\mathsf{d}_{\Gamma}$ depends on the covariance operator $\Gamma$ of the noise, which is assumed known. Thus, $\mathcal{R}_{h,B}^{\Gamma}$ itself depends on the inherent noise level in the observations.
Due to the possibility of a nonzero $h$, the map $\mathcal{R}_{h,B}^{\Gamma}$ is affine rather than linear. Its closed-form expression is provided in [315, Sec. 2, p. 5]. The associated calculations are similar to those used to establish that the posterior distribution for a linear Gaussian Bayesian inverse problem in an infinite-dimensional setting (for both data and parameters) is well-defined [71]. However, it is important to emphasize that no Gaussian assumptions are made in the current context.
@@
The expectation is taken with respect to the joint distribution $\nu\coloneqq\operatorname{Law}(u,y)$ of the random variable $(u,y)$ under (5.33). Under minor technical assumptions, the unique global minimizer of (5.36) is $(h^{\star},B^{\star})=(\overline{u},\Sigma^{1/2})$ [315, Thm. 3.1, p. 6]. Notably, even though (5.36) depends on the full joint distribution $\nu$—and hence the forward map $\mathcal{A}$—the optimal solution only depends on the first two moments of the marginal law of $u$. The technical contributions leading to this result are substantial, as they require careful analysis to handle the infinite-dimensional noise process $\eta$ and regularized inverses of compact operators.
It is also possible to develop finite-sample estimators for the problem (5.36) if only a dataset of i.i.d. pairs $\{(u_{n},y_{n})\}_{n=1}^{N}$ from $\nu$ is available. Because the optimal Tikhonov regularizer $(h^{\star},B^{\star})$ is known to be the mean and square root covariance of the random variable $u$, a natural unsupervised strategy is first to compute the empirical mean $\widehat{u}$ and empirical covariance $\widehat{\Sigma}$ from the training data, then define the plug-in estimator $(\widehat{h},\widehat{B})\coloneqq(\widehat{u},\widehat{\Sigma}^{1/2})$. Assuming that $u$ is strongly-subgaussian, application of tools from covariance estimation [315, Thm. 4.2, p. 8, and p. 29] deliver the nonasymptotic error bound